OpenStudy (anonymous):

3xy-2x²y³+4 when x is 4 and y is 6

OpenStudy (anonymous):

Easy medal

Directrix (directrix):

3 x y - 2 x² y³ + 4 Evaluate for these: x is 4 and y is 6

OpenStudy (anonymous):

For 3xy I got 72

Directrix (directrix):

3* 4* 6 - 2* (4)² * (6)³ + 4 = @usernamenotfound

OpenStudy (anonymous):

For 3xy I got 72

Directrix (directrix):

Okay.

Directrix (directrix):

What about this: - 2* (4)²

OpenStudy (anonymous):

72-32=40 what is 6³

OpenStudy (anonymous):

I got -32 for tgat

OpenStudy (anonymous):

40+,216+4=250

OpenStudy (anonymous):

No I mean 260

Directrix (directrix):

3* 4* 6 = 72 - 2* (4)² * (6)³ = -32* 6*6*6 = ?

Directrix (directrix):

What is this: -32* 6*6*6 = ?

OpenStudy (anonymous):

-32+216 184

OpenStudy (anonymous):

-32+216 184

Directrix (directrix):

The * is times

OpenStudy (anonymous):

IDK what 216(-32) is

Directrix (directrix):

-32* 6*6*6 = - 6912

OpenStudy (anonymous):

Thanks so it would be -6908

Directrix (directrix):

Put this in the Google search box and the answer will pop up: 216(-32)

OpenStudy (anonymous):

Thanks so it would be -6908

Directrix (directrix):

Not what I got. Let me check.

OpenStudy (anonymous):

Oh

Directrix (directrix):

I got -6836

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Ahh

Directrix (directrix):

Let's check

OpenStudy (anonymous):

Ahh

OpenStudy (anonymous):

Ahh

Directrix (directrix):

3* 4* 6 - 2* (4)² * (6)³ + 4 = 72 - 6912 + 4 =

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

Good job on question next

Directrix (directrix):

Thanks.